Use an appropriate local linear approximation to estimate 3√9


1
Expert's answer
2021-09-27T18:17:43-0400

The linear approximation to "f(x)=\sqrt[3]{x}"  at  "x_{0}=9" .

A linear approximation is given by  "L(x) \approx f\left(x_{0}\right)+f^{\prime}\left(x_{0}\right)\left(x-x_{0}\right) ."

We are given that  "x_{0}=9" .

Firstly, find the value of the function at the given point:  "y_{0}=f\left(x_{0}\right)=3^{\frac{2}{3}}" .

Secondly, find the derivative of the function, evaluated at the point:  "f^{\prime}(9)" .

Find the derivative:  "f^{\prime}(x)=\frac{1}{3 x^{\frac{2}{3}}}"  

Next, evaluate the derivative at the given point to find slope.

 

"f^{\prime}(9)=\frac{3^{\frac{2}{3}}}{27}"  

Plugging the values found, we get that  "L(x) \approx 3^{\frac{2}{3}}+\frac{3^{\frac{2}{3}}}{27}(x-(9))" .

Or, more simply: "L(x) \approx \frac{3^{\frac{2}{3}}}{27} x+\frac{2 \cdot 3^{\frac{2}{3}}}{3}" .

Answer:  "L(x) \approx \frac{3^{\frac{2}{3}}}{27} x+\frac{2 \cdot 3^{\frac{2}{3}}}{3} \approx 0.077040141594515 x+1.38672254870127" .